# TyphoonMLA: A Mixed Naive-Absorb MLA Kernel For Shared Prefix

### v1

@@ Abstract
Multi-Head Latent Attention (MLA) is a recent attention mechanism adopted in state-of-the-art LLMs such as DeepSeek-v3 and Kimi K2. Thanks to its novel formulation, MLA allows two functionally equivalent but computationally distinct kernel implementations: naive and absorb. While the naive kernels (e.g., FlashAttention) are typically preferred in training and prefill for their computational efficiency, existing decoding kernels (e.g., FlashMLA) rely on the absorb method to minimize HBM bandwidth usage. However, the compute-bound nature of the absorb implementations prohibits performance benefits from data reuse opportunities in attention calculations, such as shared prefixes. In this work, we introduce TyphoonMLA, a hybrid approach that combines naive and absorb formulations to harness the strengths of both. TyphoonMLA effectively leverages the shared prefix by applying the naive formulation to the compute-bound parts of attention calculations, while reducing the bandwidth requirements for non-shared parts by using the absorb formulation. As a result, TyphoonMLA improves the throughput of attention calculations in MLA architectures by up to 3× and 3.24× on NPU and GPUs, with only a 3% overhead in HBM size.
@@ 1 Introduction
In training and prefill, where performance is typically limited by the computational capacity (e.g., the total throughput of Tensor cores in GPUs), the compute-efficient naive implementation is preferred. In decode, where the HBM bandwidth typically dictates the performance, the memory-efficient absorb implementation is used. As a result, MLA architectures utilize computational resources more effectively both in training and inference, making them a more efficient alternative to other attention architectures, such as Multi-Head Attention (MHA) (Vaswani et al., 2017) or Grouped-Query Attention (GQA) (Ainslie et al., 2023). Consequently, MLA serves as the backbone of various state-of-the-art LLMs, such as DeepSeek-v3 (DeepSeek-AI et al., 2024b) and Kimi K2 (Bai et al., 2025).
@@
  • We introduce TyphoonMLA, a novel MLA inference method that achieves higher computational efficiency and resource utilization in attention calculations while producing exact outputs identical to the standard MLA kernels. To the best of our knowledge, this is the first work that combines both naive and absorb implementations in MLA calculations.
@@
Our code is open-sourced and publicly available [^1] https://github.com/acyuzuguler/TyphoonMLA-community. Generative AI tools were used to edit and refine this paper to improve the clarity and quality of writing.
@@ 2.1 Multi-Head Latent Attention
MLA is an attention mechanism introduced by DeepSeek-v2 and adopted by state-of-the-art LLMs such as DeepSeek-v3 (DeepSeek-AI et al., 2024b), Kimi K2 (Bai et al., 2025), and LongCat (Meituan et al., 2025). Unlike standard attention mechanisms, MLA employs a low-rank key-value joint compression mechanism to reduce KV-cache memory size and bandwidth requirements during inference. Moreover, its novel positional encoding enables the key-value projections to be rearranged in a way that leads to two computationally distinct implementations, namely naive and absorb.
Naive.
The naive implementation of MLA keeps the KV-cache in an uncompressed form. Fig.1(a) illustrates the components of MLA in a naive implementation. Therein, after passing through the rotational embedding and RMS normalization layers, the key and value tensors are decompressed using the up-projection matrix, denoted as $W_{KVb}$ in the figure. As a result, the K and V caches consist of multiple attention heads, and the self-attention calculation becomes equivalent to the standard MHA formulation.
Absorb.
The absorb implementation of MLA utilizes the commutative property of matrix multiplication to reposition the up-scaling matrix. In absorb formulation, as shown in Fig.1(b), the up-projection matrix is split into two submatrices (denoted as $W_{KVb1}$ and $W_{KVb2}$) and placed before and after the query ($W_{Qb}$) and output ($W_{O}$) projection layers. This formulation allows the KV-cache to remain in a compressed form, represented as PE and noPE caches in the figure. However, applying the up-projection matrix to the queries further expands their embedding dimensions, requiring more floating-point operations in the self-attention calculations. Moreover, the resulting formulation is incompatible with existing self-attention kernels (e.g., FlashAttention (Dao et al., 2022)), necessitating the development of specialized kernels for MLA layers (e.g., FlashMLA (Jiashi Li, 2025), ThunderMLA (Spector et al., 2025)).
@@ 3.1 TyphoonMLA
As previously discussed, the shared prefix introduces data reuse in the shared parts of the attention calculations. Building on this insight, TyphoonMLA partitions the attention calculation into low- and high-operational-density components and applies absorb and naive formulations, respectively. Fig.1(c) illustrates how TyphoonMLA works.
Prefill.
@@
In the prefill stage, a tree-like query structure is formed from the incoming user requests and a shared prefix (e.g., a system prompt). The queries are processed by the LLM using a prefix-aware naive kernel, producing PE and noPE caches in the low-rank latent space. The portion of the PE and noPE caches that corresponds to the shared prefix is then expanded via the up-scaling projection to form the K and V caches. Consequently, the shared and non-shared parts of the KV-cache are stored in uncompressed and compressed formats, respectively, enabling the use of both naive and absorb formulations. Notably, the up-projection incurs no additional computational overhead, as it is already performed by standard naive kernels during the prefill stage.
Decode.
@@
In the decode stage, TyphoonMLA first applies the down-projection layers $W_{Qa}$ and $W_{KVa}$ to the queries, obtaining the q and kv vectors. The q vectors pass through an RMS normalization layer, followed by the up-projection layer $W_{Qb}$. The kv vectors are split into two subvectors, one of which passes through a rotational embedding layer, and the other through an RMS normalization. All operations up to this point are common to TyphoonMLA, naive, and absorb implementations.
@@
Unlike naive and absorb implementations, TyphoonMLA performs self-attention calculations using both compressed and uncompressed KV-cache while reusing the shared operations. Similar to the naive implementation, the up-projected queries are multiplied by the uncompressed K cache, passed through a softmax function, and multiplied by the V cache. Similar to the absorb implementation, the up-projected queries are multiplied by the compressed PE and noPE caches, passed through a softmax function, and multiplied again by the noPE cache. The partial results from these two paths are then aggregated with the help of the softmax denominators. Finally, the aggregated outputs are passed through the output projection to produce the final attention outputs. Despite the different formulation and order of execution, TyphoonMLA produces outputs identical to those of the naive and absorb implementations.
Fall-back to Absorb.
@@
At small batch sizes, when there is insufficient data reuse, the naive implementation can become less efficient than absorb due to higher memory access costs. To address this, TyphoonMLA automatically switches to an absorb-only kernel whenever the batch size falls below a predefined threshold. By doing so, TyphoonMLA avoids any performance penalty at small batch sizes, ensuring consistently high efficiency across a wide range of batch sizes and operational conditions.
Parallelization.
@@
TyphoonMLA is fully compatible with existing parallelization strategies for attention computation. Although the compressed PE and noPE caches have a single head dimension, the uncompressed K and V caches can be parallelized across attention heads. Additionally, both compressed and uncompressed caches can be easily parallelized across the sequence dimension. As a result, TyphoonMLA seamlessly supports both tensor and sequence parallelism, enabling efficient scaling across multiple NPUs or GPUs.
@@
Table 1: Computational analysis of naive, absorb, and TyphoonMLA. $B$: batch size, $S_{q}$: query sequence length, $L_{s}$: shared context length, $L_{n}$: non-shared context length, $H$: number of heads, $D_{qk}$: head dim. for Q and K, $D_{v}$: head dim for V, $D_{l}$: KV LoRA rank, $D_{r}$: RoPE head dim. TyphoonMLA always requires smaller HBM reads than naive (marked in red) and fewer multiply-accumulate operations (MACs) than absorb (marked in blue).
Naive
DeepSeek-v3
MACs $BS_{q}(L_{s}+L_{n})H(D_{qk}+D_{v})$ $BL_{s}40\text{k}+BL_{n}40\text{k}$
HBM Reads $(L_{s}+BL_{n})H(D_{qk}+D_{v})$ $L_{s}40\text{k}$ ++ $BL_{n}40\text{k}$
Absorb
DeepSeek-v3
MACs $BS_{q}(L_{s}+L_{n})H(2D_{l}+D_{r})$ $BL_{s}136\text{k}$ ++ $BL_{n}136\text{k}$
HBM Reads $(L_{s}+BL_{n})(D_{l}+D_{r})$ $L_{s}576+BL_{n}576$
TyphoonMLA
DeepSeek-v3
MACs $BS_{q}L_{s}H(D_{qk}+D_{v})+BS_{q}L_{n}H(2D_{l}+D_{r})$ $BL_{s}40\text{k}$ $BL_{n}136\text{k}$
HBM Reads $L_{s}H(D_{qk}+D_{v})+BL_{n}(D_{l}+D_{r})$ $L_{s}40\text{k}$ ++ $BL_{n}576$
@@ 3.2 Computational Analysis
To quantify the potential performance benefits of the proposed method, we derive the computational requirements of attention calculations in terms of the number of multiply-accumulate operations (MACs) and HBM reads size for TyphoonMLA, as well as the naive and absorb formulations. Table 1 summarizes these derivations with respect to architectural and generation parameters. In the right-most column, we substitute the architectural parameters with those of the Deepseek-v3 to facilitate a direct comparison. In our analysis, we consider only self-attention computations, excluding the projection layers, since self-attention dominates execution time at long sequence lengths and asymptotically approaches the total runtime.
@@
Since the absorb implementation is typically compute-bound, its performance is primarily determined by the number of MACs. In contrast, the naive implementation is generally memory-bound, so its efficiency is dominated by the amount of data read from HBM. Therefore, to maximize overall performance, TyphoonMLA aims to achieve lower MACs than absorb in compute-bound regions and fewer HBM reads than naive in memory-bound regions.
@@
Table 1 shows that the naive implementation requires reading $L_{s}40\text{k}+BL_{n}40\text{k}$ numbers from HBM. In comparison, TyphoonMLA reads only $L_{s}40\text{k}+BL_{n}576$ numbers, which is the same for the shared portion but approximately 70× smaller for the non-shared portion. The absorb implementation, typically compute-bound, requires $BL_{s}136\text{k}+BL_{n}136\text{k}$ MACs. TyphoonMLA, in contrast, requires only $BL_{s}40\text{k}+BL_{n}136\text{k}$ MACs, which is the same for the non-shared portion, but 3.4× smaller in the shared portion.
This analysis reveals that TyphoonMLA requires fewer bytes to read from HBM than the naive formulation in memory-bound regions and fewer MACs than the absorb formulation in compute-bound regions, reducing the overall computational complexity of attention calculations in MLA. These reductions in both memory bandwidth and computation suggest that TyphoonMLA can achieve substantial runtime improvements, which we evaluate empirically in the next section.
@@ 4 Experiments
Building on the complexity analysis presented in Section 3.2, we now provide empirical evidence to validate the performance benefits of TyphoonMLA. In this section, we evaluate the performance of TyphoonMLA and compare it against various baselines on NPUs and GPUs. First, we measure TyphoonMLA’s throughput on popular datasets using DeepSeek-v3 and Kimi K2 with different system prompts. Next, we examine the performance breakdown of TyphoonMLA to validate our insights. Finally, we assess TyphoonMLA’s memory footprint and compare it against the absorb baseline.
Experiments on NPUs.
We implemented TyphoonMLA for Ascend NPUs using the CANN toolkit (Huawei, 2024a). For the absorb component of TyphoonMLA, we developed a custom kernel that supports paged and variable-length KV-cache using the Ascend CATLASS library (Huawei, 2025). For the naive component of TyphoonMLA, we employed the NpuFusedInferAttentionScore function from TorchNPU (Huawei, 2024b). For the projection layers and combine logic, we used TorchNPU’s einsum and vector operators.
@@
To evaluate the impact of system prompt length on performance, we selected three system prompts with varying lengths, summarized in Table 2. As benchmark datasets, we used the MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and SimpleQA (Wei et al., 2024), which comprise questions and answers from various topics. We repeat each experiment for batch sizes of 256, 512, and 1024.
@@
Fig.3 presents the normalized throughput of TyphoonMLA and the baseline methods for DeepSeek-v3 and Kimi K2. TyphoonMLA consistently outperforms both baselines across all datasets, system prompts, and batch sizes, achieving speedups between 1.2× and 3×. As expected, TyphoonMLA offers the highest speedup with Prompt A, since longer system prompts increase the ratio of shared to non-shared portions of the KV-cache. We further observe that the speedup is generally higher for the Kimi K2 than for DeepSeek-v3. This is because the number of attention heads in Kimi K2 is 64, half of DeepSeek-v3, which makes the former’s performance more sensitive to data reuse. Overall, these results confirm that TyphoonMLA effectively exploits the shared prefix and delivers consistent and significant speedups over the existing MLA kernels.
Experiments on GPUs.
@@
We also implemented TyphoonMLA using the FlashInfer naive and absorb MLA kernels (Ye et al., 2025) and repeated our NPU experiments on a GPU. Fig.4 reports the normalized throughput of TyphoonMLA on a GPU with a 1 PetaFLOPS/s of theoretical throughput in FP16, 3.3 TB/s of HBM bandwidth, comparing it against the FlashMLA (Jiashi Li, 2025) and FlashInfer (absorb-only) baselines at a batch size of 1024. TyphoonMLA achieves higher throughputs than both baselines, with factors ranging from 1.54× to 3.24×. Similar to the NPU results, TyphoonMLA accelerates Kimi K2 more than DeepSeek-v3, owing to its smaller number of attention heads. Moreover, it achieves the highest speedups with Prompt A, as it has the highest token count. Overall, these results demonstrate that TyphoonMLA generalizes effectively to GPUs, delivering substantial performance gains across diverse hardware platforms.
@@
Figure 4: Benchmark results on GPU for batch size of 1024. Y-axes represent throughput in terms of the number of generated tokens per second per layer.
Latency Breakdown.
@@
To demonstrate how TyphoonMLA accelerates attention computations and how its components behave under different batch sizes, we profiled its execution using the CANN toolkit’s msprof tool and compared it against the CATLASS absorb-only baseline. Fig.5 shows the execution time breakdown of TyphoonMLA (bars on the left-hand side) and the CATLASS baseline (bars on the right-hand side). In the figure, we denote the naive and absorb parts of attention computation in TyphoonMLA as Stage 1, Attn and Stage 2 Attn, respectively; the up and down-scaling required in the absorb formulation as $W_{KV,B1}$-proj and $W_{KV,B2}$-proj; and the combination logic that merges the intermediate results of both stages as CombineLSE. We excluded the other linear layers of the attention block, since they are identical across both methods and negligible at long sequence lengths. In this experiment, we set the shared prefix length to 4096, the non-shared sequence length of each request to 512, and used the architectural parameters of Kimi K2. Batch sizes smaller than 128 are omitted, as in these cases, TyphoonMLA falls back to the absorb-only implementation, becoming identical to the baseline. As the CATLASS baseline is absorb-only, it contains only Stage 2 components.

### v2

@@ Abstract
Multi-Head Latent Attention (MLA) is a recent attention mechanism adopted in state-of-the-art LLMs such as DeepSeek-v3 and Kimi K2. Thanks to its novel formulation, MLA allows two functionally equivalent but computationally distinct kernel implementations: naive and absorb. While the naive kernels (e.g., FlashAttention) are typically preferred in training and prefill for their computational efficiency, existing decoding kernels (e.g., FlashMLA) rely on the absorb method to minimize HBM bandwidth usage. However, the compute-bound nature of the absorb implementations prohibits performance benefits from data reuse opportunities in attention calculations, such as shared prefixes. In this work, we introduce TyphoonMLA, a hybrid approach that combines naive and absorb formulations to harness the strengths of both. TyphoonMLA effectively leverages the shared prefix by applying the naive formulation to the compute-bound parts of attention calculations, while reducing the bandwidth requirements for non-shared parts by using the absorb formulation. As a result, TyphoonMLA improves the throughput of attention calculations in MLA architectures by up to 3× and 3.24× on NPU and GPUs, and boosts end-to-end throughput by up to 1.48× in tokens per second, with only a 3% overhead in HBM size.
@@ 1 Introduction
In training and prefill, where performance is typically limited by the compute capacity (e.g., the total throughput of Tensor cores in GPUs), the compute-efficient naive implementation is preferred. In decode, where the HBM bandwidth dictates the performance, the memory-efficient absorb implementation is used. As a result, MLA architectures utilize computational resources more effectively both in training and inference, making them a more efficient alternative to other attention architectures, such as Multi-Head Attention (MHA) (Vaswani et al., 2017) or Grouped-Query Attention (GQA) (Ainslie et al., 2023). Consequently, MLA serves as the backbone of various state-of-the-art LLMs, such as DeepSeek-v3 (DeepSeek-AI et al., 2024b) and Kimi K2 (Bai et al., 2025).
@@
  • We introduce TyphoonMLA, a novel MLA inference method that achieves higher computational efficiency and resource utilization in attention calculations. To the best of our knowledge, this is the first work that combines both naive and absorb implementations in MLA calculations.
@@
Our code is open-sourced and publicly available [^1] https://github.com/huawei-csl/TyphoonMLA-community. Generative AI tools were used to edit and refine this paper to improve the clarity and quality of writing.
@@ 2.1 Multi-Head Latent Attention
MLA is an attention mechanism introduced by DeepSeek-v2 and adopted by state-of-the-art LLMs such as DeepSeek-v3 (DeepSeek-AI et al., 2024b), Kimi K2 (Bai et al., 2025), and LongCat (Meituan et al., 2025). Unlike standard attention mechanisms, MLA employs a low-rank key-value joint compression mechanism to reduce KV-cache size and bandwidth requirements during inference. Moreover, its novel positional encoding enables the key-value projections to be rearranged in a way that leads to two computationally distinct implementations, namely naive and absorb.
Naive: The naive implementation of MLA keeps the KV-cache in an uncompressed form. Fig. 1(a) illustrates the components of MLA in a naive implementation. Therein, after passing through the rotational embedding and RMS normalization layers, the key and value tensors are decompressed using the up-projection matrix, denoted as $W_{KVb}$ in the figure. As a result, the K and V caches consist of multiple attention heads, and the self-attention calculation becomes equivalent to the standard MHA formulation.
Absorb: The absorb implementation of MLA utilizes the commutative property of matrix multiplication to reposition the up-scaling matrix. In absorb formulation, as shown in Fig. 1(b), the up-projection matrix is split into two submatrices (denoted as $W_{KVb1}$ and $W_{KVb2}$) and placed before and after the query ($W_{Qb}$) and output ($W_{O}$) projection layers. This formulation allows the KV-cache to remain in a compressed form, represented as PE and noPE caches in the figure. However, applying the up-projection matrix to the queries further expands their embedding dimensions, requiring more floating-point operations in the self-attention calculations. Moreover, the resulting formulation is incompatible with existing self-attention kernels (e.g., FlashAttention (Dao et al., 2022)), necessitating the development of specialized kernels for MLA layers (e.g., FlashMLA (Jiashi Li, 2025), ThunderMLA (Spector et al., 2025)).
@@ 3.1 TyphoonMLA
As previously discussed, the shared prefix introduces data reuse in the shared parts of the attention calculations. Building on this insight, TyphoonMLA partitions the attention calculation into low- and high-arithmetic-intensity components and applies absorb and naive formulations, respectively. Fig. 1(c) illustrates how TyphoonMLA works.
@@
Prefill: In the prefill stage, a tree-like query structure is formed from the incoming user requests and a shared prefix (e.g., a system prompt). The queries are processed by the LLM using a prefix-aware naive kernel, producing PE and noPE caches in the low-rank latent space. The portion of the PE and noPE caches that corresponds to the shared prefix is then expanded via the up-projection to form the K and V caches. Consequently, the shared and non-shared parts of the KV-cache are stored in uncompressed and compressed formats, respectively, enabling the use of both naive and absorb formulations. Notably, the up-projection incurs no additional computational overhead, as it is already performed by standard naive kernels during the prefill stage.
Algorithm 1 TyphoonMLA Decode Attention Kernel
1:Queries after $\mathbf{W}_{Qb}$ proj. and RoPE layers, $\mathbf{Q}\in\mathbb{R}^{B\times H\times D_{qk}}$
2:Shared KV-cache in uncompressed form, $\mathbf{C}_{K}\in\mathbb{R}^{B\times L_{s}\times D_{qk}}$, $\mathbf{C}_{V}\in\mathbb{R}^{B\times L_{s}\times D_{v}}$
3:Non-shared KV-cache in latent space, $\mathbf{C}_{N}\in\mathbb{R}^{B\times L_{n}\times D_{l}}$, $\mathbf{C}_{R}\in\mathbb{R}^{B\times L_{n}\times D_{r}}$
4:KV up-projection matrices, $\mathbf{W}_{KVb1}\in\mathbb{R}^{H\times D_{n}\times D_{l}}$, $\mathbf{W}_{KVb2}\in\mathbb{R}^{H\times D_{v}\times D_{l}}$
5:
[…]
14:return $\mathbf{O}$
@@
Decode: In the decode stage, TyphoonMLA first applies the down-projection layers $W_{Qa}$ and $W_{KVa}$ to the queries, obtaining the q and kv vectors. The q vectors pass through an RMS normalization layer, followed by the up-projection layer $W_{Qb}$. The kv vectors are split into two subvectors, one of which passes through a rotational embedding layer, and the other through an RMS normalization. All operations up to this point are common to TyphoonMLA, naive, and absorb implementations.
@@
Unlike naive and absorb implementations, TyphoonMLA performs self-attention calculations using both compressed and uncompressed KV-cache while reusing the shared operations. As described in Algorithm 1, TyphoonMLA takes the output of $W_{Qb}$ up-projection layer ($Q$), shared KV-cache in uncompressed form ($C_{K}$ and $C_{V}$), non-shared KV-cache in latent form ($C_{N}$ and $C_{R}$), and KV up-projection matrices ($W_{KVb1}$ and $W_{KVb2}$) as inputs. The input queries $Q$ are first split from their embedding dimension into two tensors, and a positional encoding is applied. Then, similar to the naive implementation, the queries are multiplied by the uncompressed K cache, $C_{K}$, passed through a softmax function, and multiplied by the V cache, $C_{V}$. Similar to the absorb implementation, the queries are first up-projected using the $W_{KVb1}$ matrix, multiplied by the compressed PE and noPE caches ($C_{R}$ and $C_{N}$), passed through a softmax function, and multiplied again by the noPE cache. As required by the absorb formulation, the output of the absorb component is then up-projected once again using the $W_{KVb2}$ matrix. Finally, the partial results from these two components are then aggregated using a combine function with the help of the log-sum-exp (LSE) of the softmax denominators ($l_{N}$ and $l_{A}$). TyphoonMLA is mathematically equivalent to both absorb and naive implementations; therefore, it does not require any re-training or fine-tuning.
@@
Fall-back to Absorb: At small batch sizes, when there is insufficient data reuse, the naive implementation can become less efficient than absorb due to higher memory access costs. To address this, TyphoonMLA automatically switches to an absorb-only kernel whenever the batch size falls below a predefined threshold. By doing so, TyphoonMLA avoids any performance penalty at small batch sizes, ensuring consistently high efficiency across a wide range of batch sizes.
@@
Parallelization: TyphoonMLA is fully compatible with existing parallelization strategies for attention computation. Although the compressed PE and noPE caches have a single head dimension, the uncompressed K and V caches can be parallelized across attention heads. Additionally, both compressed and uncompressed caches can be easily parallelized across the sequence dimension. As a result, TyphoonMLA seamlessly supports both tensor and sequence parallelism, enabling efficient scaling across multiple NPUs or GPUs.
@@
Table 1: Computational analysis of naive, absorb, and TyphoonMLA. $B$: batch size, $S_{q}$: query sequence length, $L_{s}$: shared context length, $L_{n}$: non-shared context length, $H$: number of heads, $D_{qk}$: head dim. for Q and K, $D_{v}$: head dim for V, $D_{l}$: KV LoRA rank, $D_{n}$: noPE head dim, $D_{r}$: RoPE head dim. TyphoonMLA always requires smaller memory operations than naive (highlighted in red) and fewer multiply-accumulate operations (MACs) than absorb (highlighted in blue).
MAC DeepSeek-v3 ($\times 1024$)
Naive $BS_{q}L_{s}H(D_{qk}+D_{v})+BS_{q}L_{n}H(D_{qk}+D_{v})$ $40\times BL_{s}+40\times BL_{n}$
Absorb $BS_{q}L_{s}H(2D_{l}+D_{r})+BS_{q}L_{n}H(2D_{l}+D_{r})$ $136\times BL_{s}$ ++ $136\times BL_{n}$
TyphoonMLA $BS_{q}L_{s}H(D_{qk}+D_{v})+BS_{q}L_{n}H(2D_{l}+D_{r})$ $40\times BL_{s}$ $136\times BL_{n}$
HBM R/W DeepSeek-v3 ($\times 1024$)
Naive $L_{s}H(D_{qk}+D_{v})+BL_{n}H(D_{qk}+D_{v})$ $40\times L_{s}$ ++ $40\times BL_{n}$
Absorb $L_{s}(D_{l}+D_{r})+BL_{n}(D_{l}+D_{r})$ $0.56\times L_{s}+0.56\times BL_{n}$
TyphoonMLA $L_{s}H(D_{qk}+D_{v})+BL_{n}(D_{l}+D_{r})$ $40\times L_{s}$ ++ $0.56\times BL_{n}$
@@ 3.2 Computational Analysis
To quantify the potential performance benefits of the proposed method, we derive the computational requirements of attention calculations in terms of the number of multiply-accumulate operations (MAC) and the size of memory read/write operations (HBM R/W) for TyphoonMLA, as well as the naive and absorb formulations. Table 1 summarizes these derivations with respect to architectural and generation parameters. In the right-most column, we substitute the architectural parameters with those of the DeepSeek-v3 to facilitate a direct comparison. In our analysis, we consider only self-attention computations, excluding the projection layers, since self-attention dominates execution time at long sequence lengths and asymptotically approaches the total runtime.
@@
Since the absorb implementation is typically compute-bound, its performance is primarily determined by the number of MACs. In contrast, the naive implementation is generally memory-bound, so its efficiency is dominated by the amount of data read from HBM. Therefore, to maximize overall performance, TyphoonMLA aims to achieve lower MACs than absorb in compute-bound regions and fewer HBM read and writes than naive in memory-bound regions.
@@
Table 1 shows that the naive implementation requires reading $(40\times L_{s}+40\times BL_{n})\times 1024$ words from HBM. In comparison, TyphoonMLA reads only $(40\times L_{s}+0.56\times BL_{n})\times 1024$ words, which is the same for the shared portion but approximately 70× smaller for the non-shared portion. The absorb implementation, typically compute-bound, requires $(136\times BL_{s}+136\times BL_{n})\times 1024$ MAC operations. TyphoonMLA, in contrast, requires only $(40\times BL_{s}+136\times BL_{n})\times 1024$ MAC operations, which is the same for the non-shared portion, but 3.4× smaller in the shared portion. This analysis reveals that TyphoonMLA requires fewer bytes to read from HBM than the naive formulation in memory-bound regions and fewer MACs than the absorb formulation in compute-bound regions, reducing the overall computational complexity of attention calculations in MLA.
To combine the partial results of the naive and absorb parts, TyphoonMLA uses a CombineLSE function, similar to the epilogue stage in Flash Attention (Dao et al., 2022). This function performs only vector operations and requires reading $2BS_{q}HD_{v}$ bytes from HBM and performing $2BS_{q}HD_{v}$ MAC operations. Since the computational complexity of this function is independent of the KV sequence length, which typically ranges from hundreds to thousands, its computational overhead is insignificant compared to the other components of the attention calculations.
For TyphoonMLA to achieve a speedup, reading the shared portion of the KV-cache should be faster with the naive implementation than computation with the absorb implementation. This happens only when there is sufficient data reuse, which occurs when the batch size is larger than a threshold. By equating the memory read time of the naive and the computation time of the absorb implementation for the shared part of the KV-cache, we identify the batch size threshold, $B_{\theta}$, in terms of architectural parameters (defined above) and hardware specifications, namely throughput $T$ and HBM memory bandwidth $M$, as follows:
$\frac{L_{s}H(D_{qk}+D_{v})}{M}=\frac{{B_{\theta}}S_{q}L_{s}H(2D_{l}+D_{r})}{T}\implies B_{\theta}=\frac{(D_{qk}+D_{v})}{S_{q}(2D_{l}+D_{r})}\frac{T}{M}$ (1)
When the architectural parameters are replaced with those of DeepSeek-v3 and the hardware parameters with Ascend NPU ($T=376$ TOPS/s, $M=1.8$ TB/s), we obtain $B_{\theta}=61$. This threshold indicates the break-even point, beyond which TyphoonMLA becomes computationally more efficient than the absorb implementation, whereas, for batch sizes smaller than $B_{\theta}$, TyphoonMLA falls back to the absorb kernel to avoid any slowdown.
@@ 4 Experiments
Building on the complexity analysis presented in Section 3.2, we now provide empirical evidence to validate the performance benefits of TyphoonMLA. In this section, we evaluate the performance of TyphoonMLA and compare it against various baselines on NPUs and GPUs. First, we measure TyphoonMLA’s throughput on popular datasets using DeepSeek-v3 and Kimi K2 with different system prompts. Next, we examine its performance breakdown to validate our insights. Finally, we assess TyphoonMLA’s memory footprint and compare it against the absorb baseline.
Experiments on NPUs: We implemented TyphoonMLA for Ascend NPUs using the CANN toolkit (Huawei, 2024b). For the absorb component of TyphoonMLA, we developed a custom kernel that supports paged and variable-length KV-cache using the Ascend CATLASS library (Huawei, 2025). For the naive component of TyphoonMLA, we employed the NpuFusedInferAttentionScore function from TorchNPU (Huawei, 2024a). For the projection layers and combine logic, we used TorchNPU’s einsum and vector operators.
To evaluate the impact of system prompt length on performance, we selected three system prompts with varying lengths, summarized in Table 2. As benchmark datasets, we used the MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and SimpleQA (Wei et al., 2024), which comprise questions and answers from various topics. We repeat each experiment for batch sizes of 64, 128, 256, 512, and 1024.
@@
Fig. 2 presents the normalized throughput of TyphoonMLA and the baseline methods for the attention layers of DeepSeek-v3 and Kimi K2. TyphoonMLA consistently outperforms both baselines across all datasets, system prompts, and batch sizes, achieving speedups between 1.2× and 3×. As expected, TyphoonMLA offers the highest speedup with Prompt A, since longer system prompts increase the ratio of shared to non-shared portions of the KV-cache. We further observe that the speedup is generally higher for the Kimi K2 than for DeepSeek-v3. This is because the number of attention heads in Kimi K2 is 64, half of DeepSeek-v3, which makes the former’s performance more sensitive to data reuse. Overall, these results confirm that TyphoonMLA effectively exploits the shared prefix and delivers consistent and significant speedups over the existing MLA kernels.
@@
Experiments on GPUs: We also implemented TyphoonMLA using the FlashInfer naive and absorb MLA kernels (Ye et al., 2025) and repeated our NPU experiments on a GPU. Fig. 3 reports the normalized throughput of TyphoonMLA on a GPU with a 1 PetaFLOPS/s of theoretical throughput in FP16, 3.3 TB/s of HBM bandwidth, comparing it against the FlashMLA (Jiashi Li, 2025) and FlashInfer (absorb-only) baselines at batch sizes of 64, 128, 256, 512, and 1024. TyphoonMLA achieves higher throughputs than both baselines, with factors up to 3.24×. Similar to the NPU results, TyphoonMLA accelerates Kimi K2 more than DeepSeek-v3, owing to its smaller number of attention heads. Moreover, it achieves the highest speedups with Prompt A, as it has the highest token count. Overall, these results demonstrate that TyphoonMLA generalizes effectively to GPUs, delivering substantial performance gains across diverse hardware platforms.
@@
Figure 3: Benchmark results on GPU for various batch sizes. Y-axes represent throughput in terms of the number of generated tokens per second per layer.
Table 3: Token generation rate for DeepSeek-v3 processing MMLU dataset with a batch size of 128 per GPU.
FlashMLA TyphoonMLA
Attention time (ms) Total time (ms) TGR (kToken/s) Attention time (ms) Total time (ms) TGR (kToken/s)
Prompt A 99.1 127.2 1.01 58.1 86.3 1.48
Prompt B 34.5 62.6 2.04 25.9 54.0 2.37
Prompt C 26.9 55.0 2.33 22.0 50.1 2.56
@@
Latency Breakdown: To demonstrate how TyphoonMLA accelerates attention computations and how its components behave under different batch sizes, we profiled its execution using the CANN toolkit’s msprof tool and compared it against the CATLASS absorb-only baseline. Fig. 4 shows the execution time breakdown of TyphoonMLA (bars on the left-hand side) and the CATLASS baseline (bars on the right-hand side). In the figure, we denote the naive and absorb parts of attention computation in TyphoonMLA as Stage 1, Attn and Stage 2 Attn, respectively; the up and down-scaling required in the absorb formulation as $W_{KV,B1}$-proj and $W_{KV,B2}$-proj; and the combination logic that merges the intermediate results of both stages as CombineLSE. We excluded the other linear layers of the attention block, since they are identical across both methods and negligible at long sequence lengths. In this experiment, we set the shared prefix length to 4096, the non-shared sequence length of each request to 512, and used the architectural parameters of Kimi K2. Batch sizes smaller than 128 are omitted, as in these cases, TyphoonMLA falls back to the absorb-only implementation, becoming identical to the baseline. As the CATLASS baseline is absorb-only, it contains only Stage 2 components.